OpenStudy (anonymous):

Find the equation of the line that passes through the points (5, –3), (9, –3) in the point–slope form. A. y + 9 = 2(x - 3) B. y - 3 = 2(x + 9) C. y + 3 = 0(x - 5) D. y - 9 = -3(x + 5)

OpenStudy (anonymous):

first, get the slope by:\[\frac{ y _{1} - y _{2} }{ x _{1} - x _{2} } = m\]Where "m" is the slope. That's step #1.

OpenStudy (anonymous):

The second step is to use the point-slope formula:\[y - y _{1} = m(x - x _{1})\] where you could use either point for (x1, y1) here.

OpenStudy (anonymous):

i got -11.4 for the slope?

OpenStudy (anonymous):

woops, actually i dont know how i got -11.4 but i did it again and i got 0.

OpenStudy (anonymous):

so then would that make it c?

OpenStudy (anonymous):

@tcarroll010 ?